OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1

OpenStudy (nali):

i do not get ur question

OpenStudy (nali):

do u mean just solving for y

OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1

OpenStudy (nali):

that is the same question from above

OpenStudy (nali):

i still do not get it

OpenStudy (nali):

do u want to solve for y

OpenStudy (nali):

5x + 4y = 1 4x + 3y = -1 5x + 4y = 1 -1(4x + 3y) = -1(-1) 5x + 4y = 1 -4x + -3y = 1 add both equations --------------- x + y = 2

OpenStudy (nali):

wait i still did nto finish

OpenStudy (nali):

sorry to confuse u i have another an d easier way but just give me a few minutes to write it down

OpenStudy (nali):

solve for x in one equation and substitute in the other equation 5x + 4y = 1 4x + 3y = -1 5x + 4y =1 5x = 4y - 1 x = 4y/5 - 1/5 substitute 4x + 3y = -1 4(4y/5 - 1/5) + 3y = -1 16y/5 - 4/5 + 3y = -1 16y/5 + 3y = -1 - 4/5 31y/ 5 = -9/5

OpenStudy (nali):

 here it is